OpenStudy (anonymous):

Use elimination to find the solution to the system of equations. 4x+y=9 3x-2y=15 A. x= 13/3,y =1 B. x=9,y=-27 C. x=-1,y=13 D. x=3,y=-3

OpenStudy (anonymous):

4x+y=9 3x-2y=15 --------- multiply the top equation by 2 8x+2y = 18 3x -2y = 15 ------------ add vertically (8x+3x) + (2y-2y) = (18+15) 11x + 0 = 33 11x = 33 x = 33/11 = 3 plug x = 3 into either of the two original equations to find y

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

your welcome